# Visual Web Agents with Realistic Tasks

### v1

@@ 4.1 Strategy and System Architecture
Naïve implementations of rollout systems are synchronous in nature, like what was implemented in test-time-interaction; digirl; webrl, where a fixed group of browser sessions is forced to advance at the same pace to perform some number of rollouts in parallel. At each step, every session loads its webpage on the CPUs, the system sends all screenshots from the group together to the GPU for one combined policy call, and then waits until the slowest session completes before any session proceeds; a common variant waits at the end of each full trajectory instead of at every step. These step and episode barriers are problematic in web settings because step times and horizons vary widely across tasks and even across data-collection policies. Faster sessions sit idle while slower ones catch up, and hardware activity arrives in synchronized bursts that saturate either CPUs or GPUs at once, and then drop to idle. We illustrate this “burst-idle” behavior in Figure 6 and then quantify its impact on end-to-end data collection time and scaling behavior in Figure 7.
@@
To mitigate the inefficiency from this “burst-idle” behavior, our framework adopts a server/client architecture (server-client-arch), where the server hosts simulation environments and the client organizes the rollout loop. The server simulates the browser environments following a master/worker paradigm (master-worker-paradigm), with a master node routing API requests to CPU worker nodes that host the actual simulators and execute actions. The master node configuration determines how much parallel work the system can support, making it easy to scale horizontally by adding workers or vertically by increasing per-node resources. The server is entirely CPU-based and requires no GPUs. The client hosts agent instances on the GPU and send requests to the server. To manage the requests sent to the server, we adapt an operation-specific local queue that enables spreading CPU and GPU workload evenly under tight resource settings, with more details in §B.1. The client is GPU-heavy and relies little on CPU budgets.
@@ 5.1 Experimental Setup
Action space. We parameterize the action space by extending the definition of test-time-interaction with a navigation action, so that the action space now includes clicking, typing, scrolling, returning to the previous page, and navigating to a website specified by the agent. The WebGym rollout system supports both the Set-of-Marks mode (set-of-marks), where interactive elements are labeled in the screenshots by the simulator, and a coordinate-based screenshot-only mode. We operate directly in the coordinate-based mode, which is supported by the base model (Qwen3-VL-8B) we use to train via RL.
Policy update. We run REINFORCE (original-reinforce) with binary terminal rewards, without any baseline or negative gradient. This update is equivalent to online filtered behavior cloning (BC) (digirl) or a “thresholded” version of reward-weighted regression (reward-weighted-regression), which retains only successful trajectories and maximizes their log-likelihood. The training objective is given by:
@@
Base policy. We use the Qwen3-VL model family for its strong grounding and reasoning capabilities (qwen3-vl; qwen3vl-repo). We experiment on the 8B-sized model with both Instruct and Thinking variants.
@@
Probing task hardness with proprietary VLMs. To understand the hardness of the OOD test set in WebGym, we also evaluate GPT-4o and GPT-5 (Thinking) as two strong proprietary models on the OOD test set. Results show that simply using the same prompt as Qwen3-VL (§E.1) with Set-of-Marks (SoM (set-of-marks), used here because these models cannot accurately output coordinates) leads to non-trivial but far below 100% performance, where GPT-4o achieves 27.1% on the full test set, and GPT-5 achieves 29.8% on a randomly sampled subset with 300 tasks from the OOD test set (subsampled due to the budget costs associated with GPT-5). Note that GPT-4o achieves 25.6% on this task subset, so the variance caused by subsampling is not significant. As we discuss next, our trained agent significantly outperforms both of these compared approaches.
@@
Figure 8: Ablations on base models, prompting, and action filtering. (left) test-set success rate curves of different models (Qwen3-VL-Instruct-8B, Qwen3-VL-Thinking-8B, GPT-4o, and GPT-5-Thinking), and of Qwen3-VL-Instruct-8B under different constraints (either removing the memory prompt or removing the repetition penalty during RL). (right) test-set same-screenshot (repetitive inefficient action) rate and trajectory length curves before and after applying the action repetition penalty (filtering) on the Instruct model. The hyper-parameters of the best run (“Instruct-8B”) are shown in §F. The orange curve is an average of two runs, while the translucent orange area around the orange line denotes the variance of these two runs. We observe that even if the starting variance between different runs can be large, the curves converges after 24k training trajectories are collected. All plots are produced with an EMA (Exponential Moving Average) smoothing of $0.50$.
@@ 5.2 Design Choices for Agent Training
In this subsection, we describe the key design choices we had to utilize to train agents via RL on WebGym. Inspired by prior research on RL for LLMs (e3; reinforce-ada), we note that biasing the sampled tasks towards hard tasks improves the performance and sample efficiency, as long as easy tasks are also used either via co-training or via a prior stage of a curriculum. Thus, we start from a difficulty sampling that slightly biases towards harder tasks (concretely, easy : medium : hard = $2:5:3$). All experiments in this section follow this task sampling strategy, while we study different sampling strategies further in §5.3.
Memory prompt for long-horizon tasks. Our base model (Qwen3-VL-8B) follows a simple prompting template for computer-use agents, which includes a high-level action description (e.g., “click on the search button”) paired with a low-level tool call (e.g., “Click [209, 441]”) (qwen3-vl; qwen3vl-repo). The exact prompt is shown in §E.1. From preliminary experiments, we observe that this template yields strong initial performance, because the Qwen-3 series of models are trained on data collected from web environments with this prompt (qwen3-vl; qwen3vl-repo), but this prompt provides limited improvement if we use it to rollout trajectories with RL on WebGym, as shown in Figure 9 (left).
This limitation arises from the model’s inability to retain and reuse information gathered in earlier steps. For example, in the task shown in Figure 2, where the goal is to compare two products, the agent must recall details about the first product when evaluating the second. In the current design, this requires including the entire conversation history and all prior screenshots as input, which is inefficient and impractical. A more formal treatment of this challenge from a partially observable Markov decision process (POMDP) perspective is provided in §C. To solve this limitation, we propose that the model should maintain a memory by responding with the updated memory in each step, inspired by GLM-4.1v (glm-4.1v). The prompt with memory is also demonstrated in §E.1. Utilizing the memory prompt significantly boosts performance after RL (see Figure 9, left, Instruct-8B versus w/o Memory Prompt).
@@ 5.3 Main Results: Scaling Agent Performance with WebGym
Table 2: WebGym empowers the small Qwen3-VL-Instruct-8B model to learn with RL to achieve state-of-the-art performance on the holdout test task set, arriving at a strikingly 42.9% success rate, significantly surpassing all agents empowered by proprietary models and small-yet-effective open-source agentic models. *The Official prompt refers to the web agent prompt specified in qwen3vl-repo.
@@
All main results from this section are summarized in Table 3. Our final agent is trained with the memory prompt and penalty for repetitive actions, with a task horizon of $(10,20,30)$ over (easy, medium, hard) tasks, and with uniform sampling across all task difficulties, surpassing the best proprietary model evaluated (GPT-5-Thinking) by 13.1%, as shown in Table 2 and Table 3. We provide raw results to promote fair comparison, and encourage readers to include raw results in future work to compare with these scores.
@@ 6 Discussion and Conclusion
Future work. Despite the availability of meaningful task-specific rubrics in our task set, LLM-generated rubrics can sometimes be overly strict, which slightly reduces sample efficiency during training. An important direction for future work is to explore more advanced designs for rubric-based evaluation. One promising approach is to treat LLM-generated rubrics as soft guidance rather than strict criteria, for example by training an evaluator that jointly conditions on both the rubric and the task instruction. Another promising direction is to investigate reinforcement learning algorithms that utilize WebGym more efficiently, particularly methods based on dynamic curricula (test-time-interaction), automatic curriculum learning (digirl), multi-agent training, and cross-site navigation.
@@ Appendix C Modeling the Markov Decision Process On Long-Horizon Web Navigation Tasks
Recent work models the web agent tasks as a POMDP, because historical observations must be included to successfully infer the next action. Formally, for a web agent task, if we represent the observation (including the screenshot and metadata of that webpage) at step $t$ as $o_{t}$, recent work represents the true state $s_{t}$ with $s_{t}=o_{t-n:t}$ (concatenation of the observations of last $n$ steps) (qwen3-vl; test-time-interaction). This is called a windowed history, designed to fit into the context limit of the vision-language models. Formally, this modeling assumes the approximation
@@
where $o$ is the vector that represents observation, and $c$ is the vector that represents task (pae-webvoyager; test-time-interaction; qwen3-vl; qwen3vl-repo). In words, these works treat observations of the last several steps as sufficient statistics of history for predicting future observations. In practice, the window size $n$ is usually set to $3$. For long-horizon tasks, this modeling over-simplifies the problem because the policy will frequently return to the same state. If the policy is not aware of its full history, it can produce repetitive sequences of actions because it is not aware of what it already finished earlier.
@@
Practically, the problem with inputting full history observations $o_{1:t}$ to the VLM policy is that the context window of even the most advanced VLMs is not large enough, partially because image is too dense a representation for web agents tasks. Recent work like GLM-4.1v (glm-4.1v) addresses this limitation by introducing a memory as part of the tokens that the VLM should output in each step. Formally, the memory at step $t$, denoted by $m_{t}$, is a result of the observation without memory $\tilde{o}_{t}$, the memory from the previous step $m_{t-1}$, and the task instruction $c$, denoted as $m_{t}=\text{VLM}_{\theta}(\tilde{o}_{t},m_{t-1},c)$, and the observation at step $t$ can now be denoted as $o_{t}=(m_{t},\tilde{o}_{t},c)$. This design leads to a continuous compression of the memory based on the progress of the task, which resembles RLHF and Long-Short-Term Memory (LSTM). We practice this idea for our agent context management, as shown in Section E.1.
@@ E.1 Agent Action Prompt
The prompt we use to rollout trajectories is the agent action prompts shown below. Both the official prompt as provided in qwen3vl-repo and the memory prompt that we use across §5 are included below.

### v2

@@ 4.1 Strategy and System Architecture
Naïve implementations of rollout systems are synchronous in nature, like what was implemented in Shen et al. (2025); Bai et al. (2024); Qi et al. (2024), where a fixed group of browser sessions is forced to advance at the same pace to perform some number of rollouts in parallel. At each step, every session loads its webpage on the CPUs, the system sends all screenshots from the group together to the GPU for one combined policy call, and then waits until the slowest session completes before any session proceeds; a common variant waits at the end of each full trajectory instead of at every step. These step and episode barriers are problematic in web settings because step times and horizons vary widely across tasks and even across data-collection policies. Faster sessions sit idle while slower ones catch up, and hardware activity arrives in synchronized bursts that saturate either CPUs or GPUs at once, and then drop to idle. We illustrate this “burst-idle” behavior in Figure 6 and then quantify its impact on end-to-end data collection time and scaling behavior in Figure 7.
@@
To mitigate the inefficiency from this “burst-idle” behavior, our framework adopts a server/client architecture (Ali et al., 2020), where the server hosts simulation environments and the client organizes the rollout loop. The server simulates the browser environments following a master/worker paradigm (Aida et al., 2003), with a master node routing API requests to CPU worker nodes that host the actual simulators and execute actions. The master node configuration determines how much parallel work the system can support, making it easy to scale horizontally by adding workers or vertically by increasing per-node resources. The server is entirely CPU-based and requires no GPUs. The client hosts agent instances on the GPU and send requests to the server. To manage the requests sent to the server, we adapt an operation-specific local queue that enables spreading CPU and GPU workload evenly under tight resource settings, with more details in §B.1. The client is GPU-heavy and relies little on CPU budgets.
@@ 5.1 Experimental Setup
Action space. We parameterize the action space by extending the definition of Shen et al. (2025) with a navigation action, so that the action space now includes clicking, typing, scrolling, returning to the previous page, and navigating to a website specified by the agent. The WebGym rollout system supports both the Set-of-Marks mode (Yang et al., 2023), where interactive elements are labeled in the screenshots by the simulator, and a coordinate-based screenshot-only mode. We operate directly in the coordinate-based mode, which is supported by the base model (Qwen3-VL-8B) we use to train via RL.
Policy update. We run REINFORCE (Williams, 1992) with binary terminal rewards, without any baseline or negative gradient. This update is equivalent to online filtered behavior cloning (BC) (Bai et al., 2024) or a “thresholded” version of reward-weighted regression (Peters and Schaal, 2007), which retains only successful trajectories and maximizes their log-likelihood. The training objective is given by:
@@
Base policy. We use the Qwen3-VL model family for its strong grounding and reasoning capabilities (Bai et al., 2025b; QwenLM-Team, 2025). We experiment on the 8B-sized model with both Instruct and Thinking variants.
@@
Probing task hardness with proprietary VLMs. To understand the hardness of the OOD test set in WebGym, we also evaluate GPT-4o and GPT-5 (Thinking) as two strong proprietary models on the OOD test set. Results show that simply using the same prompt as Qwen3-VL (§E.1) with Set-of-Marks (SoM (Yang et al., 2023), used here because these models cannot accurately output coordinates) leads to non-trivial but far below 100% performance, where GPT-4o achieves 27.1% on the full test set, and GPT-5 achieves 29.8% on a randomly sampled subset with 300 tasks from the OOD test set (subsampled due to the budget costs associated with GPT-5). Note that GPT-4o achieves 25.6% on this task subset, so the variance caused by subsampling is not significant. As we discuss next, our trained agent significantly outperforms both of these compared approaches.
@@ 5.2 Design Choices for Agent Training
In this subsection, we describe the key design choices we had to utilize to train agents via RL on WebGym. Inspired by prior research on RL for LLMs (Setlur et al., 2025; Xiong et al., 2025), we note that biasing the sampled tasks towards hard tasks improves the performance and sample efficiency, as long as easy tasks are also used either via co-training or via a prior stage of a curriculum. Thus, we start from a difficulty sampling that slightly biases towards harder tasks (concretely, easy : medium : hard = $2:5:3$). All experiments in this section follow this task sampling strategy, while we study different sampling strategies further in §5.3.
Memory prompt for long-horizon tasks. Our base model (Qwen3-VL-8B) follows a simple prompting template for computer-use agents, which includes a high-level action description (e.g., “click on the search button”) paired with a low-level tool call (e.g., “Click [209, 441]”) (Bai et al., 2025b; QwenLM-Team, 2025). The exact prompt is shown in §E.1. From preliminary experiments, we observe that this template yields strong initial performance, because the Qwen-3 series of models are trained on data collected from web environments with this prompt (Bai et al., 2025b; QwenLM-Team, 2025), but this prompt provides limited improvement if we use it to rollout trajectories with RL on WebGym, as shown in Figure 8 (left).
This limitation arises from the model’s inability to retain and reuse information gathered in earlier steps. For example, in the task shown in Figure 2, where the goal is to compare two products, the agent must recall details about the first product when evaluating the second. In the current design, this requires including the entire conversation history and all prior screenshots as input, which is inefficient and impractical. A more formal treatment of this challenge from a partially observable Markov decision process (POMDP) perspective is provided in §C. To solve this limitation, we propose that the model should maintain a memory by responding with the updated memory in each step, inspired by GLM-4.1v (Hong et al., 2025). The prompt with memory is also demonstrated in §E.1. Utilizing the memory prompt significantly boosts performance after RL (see Figure 8, left, Instruct-8B versus w/o Memory Prompt).
@@ 5.3 Main Results: Scaling Agent Performance with WebGym
Table 2: WebGym empowers the small Qwen3-VL-Instruct-8B model to learn with RL to achieve state-of-the-art performance on the holdout test task set, arriving at a strikingly 42.9% success rate, significantly surpassing all agents empowered by proprietary models and small-yet-effective open-source agentic models. *The Official prompt refers to the web agent prompt specified in QwenLM-Team (2025).
@@
Setting Legend in Figure 9 Peak Performance (%) $\Delta$ from Previous (%)
Baseline Exclude Domains 31.0
+Breadth Biased to Hard 34.5 +3.5
+Size Only Easy 36.9 +2.4
+Depth Uniform Sampling 38.2 +1.3
+Horizon Control Shorten Horizon 42.9 +4.8
Table 3: Scaling overview: peak performance of different scaling directions through RL settings on the test set. $\Delta$ from Previous shows the improvement over the immediately preceding setting.
All main results from this section are summarized in Table 4, and the overview of the scaling results we discussed is presented in Table 3. Our final agent is trained with the memory prompt and penalty for repetitive actions, with a task horizon of $(10,20,30)$ over (easy, medium, hard) tasks, and with uniform sampling across all task difficulties, surpassing the best proprietary model evaluated (GPT-5-Thinking) by 13.1%, as shown in Table 2 and Table 4. We provide raw results to promote fair comparison, and encourage readers to include raw results in future work to compare with these scores.
@@ 6 Discussion and Conclusion
Future work. Despite the availability of meaningful task-specific rubrics in our task set, LLM-generated rubrics can sometimes be overly strict, which slightly reduces sample efficiency during training. An important direction for future work is to explore more advanced designs for rubric-based evaluation. One promising approach is to treat LLM-generated rubrics as soft guidance rather than strict criteria, for example by training an evaluator that jointly conditions on both the rubric and the task instruction. Another promising direction is to investigate reinforcement learning algorithms that utilize WebGym more efficiently, particularly methods based on dynamic curricula (Shen et al., 2025), automatic curriculum learning (Bai et al., 2024), multi-agent training, and cross-site navigation.
@@ Appendix C Modeling the Markov Decision Process On Long-Horizon Web Navigation Tasks
Recent work models the web agent tasks as a POMDP, because historical observations must be included to successfully infer the next action. Formally, for a web agent task, if we represent the observation (including the screenshot and metadata of that webpage) at step $t$ as $o_{t}$, recent work represents the true state $s_{t}$ with $s_{t}=o_{t-n:t}$ (concatenation of the observations of last $n$ steps) (Bai et al., 2025b; Shen et al., 2025). This is called a windowed history, designed to fit into the context limit of the vision-language models. Formally, this modeling assumes the approximation
@@
where $o$ is the vector that represents observation, and $c$ is the vector that represents task (Zhou et al., 2024; Shen et al., 2025; Bai et al., 2025b; QwenLM-Team, 2025). In words, these works treat observations of the last several steps as sufficient statistics of history for predicting future observations. In practice, the window size $n$ is usually set to $3$. For long-horizon tasks, this modeling over-simplifies the problem because the policy will frequently return to the same state. If the policy is not aware of its full history, it can produce repetitive sequences of actions because it is not aware of what it already finished earlier.
@@
Practically, the problem with inputting full history observations $o_{1:t}$ to the VLM policy is that the context window of even the most advanced VLMs is not large enough, partially because image is too dense a representation for web agents tasks. Recent work like GLM-4.1v (Hong et al., 2025) addresses this limitation by introducing a memory as part of the tokens that the VLM should output in each step. Formally, the memory at step $t$, denoted by $m_{t}$, is a result of the observation without memory $\tilde{o}_{t}$, the memory from the previous step $m_{t-1}$, and the task instruction $c$, denoted as $m_{t}=\text{VLM}_{\theta}(\tilde{o}_{t},m_{t-1},c)$, and the observation at step $t$ can now be denoted as $o_{t}=(m_{t},\tilde{o}_{t},c)$. This design leads to a continuous compression of the memory based on the progress of the task, which resembles RLHF and Long-Short-Term Memory (LSTM). We practice this idea for our agent context management, as shown in Section E.1.
@@ E.1 Agent Action Prompt
The prompt we use to rollout trajectories is the agent action prompts shown below. Both the official prompt as provided in QwenLM-Team (2025) and the memory prompt that we use across §5 are included below.